OpenStudy (anonymous):

How do you find a complex solution? ie. x^2-18x+78=-8

OpenStudy (anonymous):

\[x^2-18x+78=-8\\ x^2-18x=-86\] is a start

OpenStudy (anonymous):

then \[(x-9)^2=-86+81=-5\]

OpenStudy (anonymous):

and so \[x-9=\pm \sqrt{5}i\]

OpenStudy (anonymous):

Okay thank you!!!!!!!!!!!!

OpenStudy (anonymous):

yw